# Complex Wavelet Mutual Information Loss: A Multi-Scale Loss Function for Semantic Segmentation

### v1

@@ 3.2 Mutual Information in Complex Wavelet Domain
CE $0.770_{\pm 0.020}$ $0.856_{\pm 0.017}$ $1.39_{\pm 0.19}$ $0.397_{\pm 0.091}$ $2.37_{\pm 0.51}$ $0.754_{\pm 0.020}$ $0.843_{\pm 0.017}$ $1.43_{\pm 0.19}$ $0.381_{\pm 0.067}$ $2.90_{\pm 1.10}$
BCE $0.742_{\pm 0.004}$ $0.835_{\pm 0.003}$ $1.53_{\pm 0.19}$ $0.372_{\pm 0.064}$ $2.50_{\pm 0.36}$ $0.750_{\pm 0.016}$ $0.841_{\pm 0.013}$ $1.44_{\pm 0.13}$ $0.417_{\pm 0.050}$ $2.21_{\pm 0.14}$
Dice $0.775_{\pm 0.027}$ $0.860_{\pm 0.020}$ $1.32_{\pm 0.17}$ $0.458_{\pm 0.055}$ $1.85_{\pm 0.14}$ $0.766_{\pm 0.009}$ $0.852_{\pm 0.006}$ $1.38_{\pm 0.19}$ $0.421_{\pm 0.117}$ $2.70_{\pm 0.89}$
Focal $0.734_{\pm 0.004}$ $0.829_{\pm 0.002}$ $1.40_{\pm 0.12}$ $0.475_{\pm 0.122}$ $2.20_{\pm 0.47}$ $0.740_{\pm 0.015}$ $0.833_{\pm 0.012}$ $1.39_{\pm 0.17}$ $0.480_{\pm 0.079}$ $2.02_{\pm 0.47}$
Jaccard $0.760_{\pm 0.011}$ $0.849_{\pm 0.010}$ $1.29_{\pm 0.27}$ $0.492_{\pm 0.138}$ $2.20_{\pm 0.94}$ $0.768_{\pm 0.013}$ $0.856_{\pm 0.010}$ $1.25_{\pm 0.15}$ $0.533_{\pm 0.069}$ $\underline{1.55_{\pm 0.25}}$
Tversky $0.754_{\pm 0.028}$ $0.845_{\pm 0.022}$ $1.29_{\pm 0.07}$ $0.526_{\pm 0.066}$ $1.54_{\pm 0.47}$ $0.772_{\pm 0.017}$ $0.859_{\pm 0.013}$ $1.26_{\pm 0.07}$ $0.523_{\pm 0.015}$ $1.65_{\pm 0.36}$
WCE $0.735_{\pm 0.014}$ $0.829_{\pm 0.012}$ $1.44_{\pm 0.10}$ $0.454_{\pm 0.019}$ $2.30_{\pm 0.70}$ $0.732_{\pm 0.012}$ $0.828_{\pm 0.010}$ $1.45_{\pm 0.03}$ $0.455_{\pm 0.045}$ $2.04_{\pm 0.51}$
ABW - - - - - - - - - -
Skea-topo - - - - - $0.669_{\pm 0.008}$ $0.772_{\pm 0.014}$ $1.57_{\pm 0.14}$ $0.397_{\pm 0.140}$ $5.79_{\pm 4.49}$
RMI $\underline{0.786_{\pm 0.015}}$ $\underline{0.868_{\pm 0.011}}$ $1.31_{\pm 0.14}$ $0.436_{\pm 0.073}$ $1.83_{\pm 0.27}$ $\mathbf{0.786_{\pm 0.002}}$ $\mathbf{0.868_{\pm 0.002}}$ $1.25_{\pm 0.08}$ $0.467_{\pm 0.024}$ $1.75_{\pm 0.20}$
clDice $0.466_{\pm 0.072}$ $0.575_{\pm 0.055}$ $2.05_{\pm 0.13}$ $0.111_{\pm 0.089}$ $14.28_{\pm 10.08}$ $0.468_{\pm 0.048}$ $0.577_{\pm 0.052}$ $1.97_{\pm 0.10}$ $0.181_{\pm 0.145}$ $17.52_{\pm 12.82}$
CWMI-Real $0.774_{\pm 0.018}$ $0.861_{\pm 0.013}$ $\underline{1.12_{\pm 0.18}}$ $\underline{0.590_{\pm 0.115}}$ $\underline{1.45_{\pm 0.31}}$ $\underline{0.782_{\pm 0.050}}$ $\underline{0.865_{\pm 0.038}}$ $\underline{1.13_{\pm 0.16}}$ $\underline{0.566_{\pm 0.038}}$ $1.59_{\pm 0.91}$
CWMI $\mathbf{0.801_{\pm 0.011}}$ $\mathbf{0.880_{\pm 0.008}}$ $\mathbf{1.06_{\pm 0.08}}$ $\mathbf{0.596_{\pm 0.032}}$ $\mathbf{1.07_{\pm 0.13}}$ $0.778_{\pm 0.026}$ $0.863_{\pm 0.019}$ $\mathbf{1.01_{\pm 0.05}}$ $\mathbf{0.641_{\pm 0.024}}$ $\mathbf{1.48_{\pm 0.47}}$
@@
CE $0.730_{\pm 0.020}$ $0.823_{\pm 0.018}$ $3.80_{\pm 0.63}$ $0.148_{\pm 0.087}$ $11.32_{\pm 4.41}$ $0.739_{\pm 0.021}$ $0.831_{\pm 0.016}$ $2.94_{\pm 0.92}$ $0.280_{\pm 0.155}$ $10.64_{\pm 0.98}$
BCE $0.684_{\pm 0.008}$ $0.790_{\pm 0.008}$ $1.66_{\pm 0.30}$ $0.575_{\pm 0.058}$ $13.01_{\pm 3.64}$ $0.691_{\pm 0.002}$ $0.796_{\pm 0.002}$ $1.63_{\pm 0.19}$ $0.577_{\pm 0.049}$ $13.00_{\pm 4.04}$
Dice $0.762_{\pm 0.013}$ $0.850_{\pm 0.011}$ $1.58_{\pm 0.15}$ $0.560_{\pm 0.049}$ $11.39_{\pm 6.13}$ $0.760_{\pm 0.010}$ $0.848_{\pm 0.009}$ $1.64_{\pm 0.20}$ $0.550_{\pm 0.038}$ $11.61_{\pm 5.15}$
Focal $0.677_{\pm 0.013}$ $0.785_{\pm 0.011}$ $1.80_{\pm 0.18}$ $0.540_{\pm 0.055}$ $11.80_{\pm 2.09}$ $0.678_{\pm 0.017}$ $0.786_{\pm 0.014}$ $1.74_{\pm 0.13}$ $0.550_{\pm 0.043}$ $12.76_{\pm 2.50}$
Jaccard $0.759_{\pm 0.009}$ $0.848_{\pm 0.007}$ $1.34_{\pm 0.03}$ $0.629_{\pm 0.005}$ $11.22_{\pm 4.14}$ $0.753_{\pm 0.007}$ $0.844_{\pm 0.006}$ $1.36_{\pm 0.12}$ $0.626_{\pm 0.016}$ $11.65_{\pm 7.52}$
Tversky $0.753_{\pm 0.010}$ $0.844_{\pm 0.008}$ $1.35_{\pm 0.25}$ $0.628_{\pm 0.078}$ $11.61_{\pm 5.02}$ $0.749_{\pm 0.003}$ $0.841_{\pm 0.003}$ $1.56_{\pm 0.43}$ $0.581_{\pm 0.101}$ $11.76_{\pm 2.12}$
WCE $0.651_{\pm 0.011}$ $0.763_{\pm 0.010}$ $1.71_{\pm 0.08}$ $0.556_{\pm 0.003}$ $\underline{10.94_{\pm 0.37}}$ $0.649_{\pm 0.008}$ $0.761_{\pm 0.007}$ $1.82_{\pm 0.14}$ $0.527_{\pm 0.046}$ $12.37_{\pm 2.06}$
ABW $0.688_{\pm 0.011}$ $0.794_{\pm 0.009}$ $1.61_{\pm 0.11}$ $0.580_{\pm 0.034}$ $11.80_{\pm 3.22}$ $0.680_{\pm 0.005}$ $0.788_{\pm 0.004}$ $1.66_{\pm 0.10}$ $0.568_{\pm 0.035}$ $\underline{10.47_{\pm 2.00}}$
Skea-topo $0.620_{\pm 0.095}$ $0.734_{\pm 0.085}$ $2.24_{\pm 0.82}$ $0.463_{\pm 0.159}$ $11.65_{\pm 5.62}$ $0.673_{\pm 0.005}$ $0.780_{\pm 0.004}$ $2.02_{\pm 0.26}$ $0.500_{\pm 0.058}$ $11.28_{\pm 1.14}$
RMI $\underline{0.764_{\pm 0.002}}$ $\underline{0.851_{\pm 0.001}}$ $1.66_{\pm 0.14}$ $0.541_{\pm 0.024}$ $11.57_{\pm 2.94}$ $\underline{0.762_{\pm 0.006}}$ $\underline{0.849_{\pm 0.003}}$ $1.92_{\pm 0.45}$ $0.474_{\pm 0.093}$ $11.68_{\pm 4.24}$
clDice $0.653_{\pm 0.060}$ $0.762_{\pm 0.053}$ $2.28_{\pm 0.63}$ $0.459_{\pm 0.103}$ $15.55_{\pm 4.59}$ $0.683_{\pm 0.038}$ $0.788_{\pm 0.031}$ $1.74_{\pm 0.21}$ $0.559_{\pm 0.034}$ $13.54_{\pm 1.88}$
CWMI-Real $0.761_{\pm 0.012}$ $0.850_{\pm 0.010}$ $\underline{1.25_{\pm 0.29}}$ $\underline{0.653_{\pm 0.069}}$ $\mathbf{10.76_{\pm 2.17}}$ $0.759_{\pm 0.012}$ $0.849_{\pm 0.009}$ $\underline{1.25_{\pm 0.10}}$ $\underline{0.647_{\pm 0.027}}$ $10.90_{\pm 3.70}$
CWMI $\mathbf{0.773_{\pm 0.007}}$ $\mathbf{0.859_{\pm 0.005}}$ $\mathbf{1.15_{\pm 0.23}}$ $\mathbf{0.664_{\pm 0.071}}$ $11.17_{\pm 7.44}$ $\mathbf{0.770_{\pm 0.011}}$ $\mathbf{0.856_{\pm 0.009}}$ $\mathbf{1.18_{\pm 0.10}}$ $\mathbf{0.657_{\pm 0.029}}$ $\mathbf{10.37_{\pm 7.02}}$
Table 1: Quantitative results of different loss functions across the four datasets and two neural network models. The bold numbers indicate the best performance for each metric, while the underlined numbers denote the second-best performance. A hyphen (”-”) indicates cases where the model did not converge.
@@ 4.1 Experimental Setup
To evaluate CWMI loss, we employed U-Net (Ronneberger et al., 2015) and Attention U-Net (Oktay et al., 2018) as base models. This choice allows us to assess CWMI’s generalization across fully convolutional models and attention-based models.
@@
We tested CWMI on three public segmentation datasets, all characterized by class and instance imbalance: (1) SNEMI3D, a neurite segmentation dataset containing 100 $1024\times 1024$ grayscale images from electron microscopy slices (Arganda-Carreras et al., 2013); (2) GlaS, a gland segmentation dataset with 165 RGB images of varying sizes from histological images of colorectal cancer samples (Sirinukunwattana et al., 2017); (3) DRIVE, a retinal vessel segmentation dataset comprising 40 $584\times 565$ RGB images from fundus photographs (Staal et al., 2004); and (4) the Massachusetts Roads dataset (MASS ROAD), a road segmentation dataset with 1171 $1500\times 1500$ RGB images from aerial imagery (Mnih, 2013). We choose a subset of 120 images (ignoring images without a network of roads). For all datasets, three-fold cross-validation was used to ensure robust evaluation.
@@
We compared CWMI against 11 state-of-the-art (SOTA) loss functions, including pixel-wise loss functions (e.g., cross entropy, BCE (Long et al., 2015), Focal loss (Ross & Dollár, 2017)), region-based loss functions (e.g., Dice loss (Milletari et al., 2016), Tversky loss (Salehi et al., 2017), Jaccard loss (Rahman & Wang, 2016)), and structural/topological loss functions (e.g., WCE (Ronneberger et al., 2015), ABW loss (Liu et al., 2022), Skea-topo loss (Liu et al., 2024), RMI loss (Zhao et al., 2019), clDice loss (Shit et al., 2021)). Hyperparameters for each baseline were tuned via grid search, with Tversky loss ($\alpha=0.5,\beta=0.5$) and Focal loss ($\gamma=2.5$) as examples.
We utilized steerable pyramids with four decomposition levels and four orientations in all experiments. A regularization parameter of $\lambda=0.1$ was applied in all CWMI experiments. For the implementation of Equation 10, although PyTorch supports complex matrix calculations, our experiments indicated that its efficiency remains suboptimal. Consequently, we computed using their real representations, which are mathematically equivalent (Golub & Van Loan, 2013). To assess the significance of the complex steerable pyramid, we compared CWMI with a real-number-only variant (CWMI-Real), implemented according to Equations 4 and 9. Adam optimizer with a StepLR scheduler (initial learning rate $1\times 10^{-4}$, decay rate 0.8, step size 10) was used, and models were trained for 50 epochs with a batch size of 10. Early stopping based on mIoU was employed to select the best model. Training was conducted on an NVIDIA A100 GPU using the Google Colab runtime.
@@
Figure 3: Qualitative results of different loss functions on the SNEMI3D dataset. Red: false positive regions; Blue: false negative regions. Green arrow: challenging false positive and Organge arrow: challenging false negative that are successfully addressed by CWMI.

### v2

@@ 3.2 Mutual Information in Complex Wavelet Domain
CE .$770_{\pm.014}.$$856_{\pm.011}.$ $1.379_{\pm.136}$$406_{\pm.059}.$ $2.344_{\pm.398}$  .$757_{\pm.017}.$$845_{\pm.014}.$ $1.436_{\pm.126}$$372_{\pm.059}.$ $2.932_{\pm.875}$
BCE .$742_{\pm.004}.$$835_{\pm.003}.$ $1.501_{\pm.175}$$399_{\pm.088}.$ $2.355_{\pm.434}$  .$746_{\pm.012}.$$838_{\pm.010}.$ $1.443_{\pm.106}$$427_{\pm.033}.$ $2.169_{\pm.264}$
Dice .$779_{\pm.018}.$$863_{\pm.014}.$ $1.293_{\pm.116}$$471_{\pm.045}.$ $1.696_{\pm.275}$  .$776_{\pm.015}.$$861_{\pm.012}.$ $1.335_{\pm.136}$$445_{\pm.079}.$ $2.146_{\pm.832}$
Focal .$737_{\pm.008}.$$832_{\pm.006}.$ $1.404_{\pm.168}$$471_{\pm.113}.$ $2.147_{\pm.331}$  .$745_{\pm.011}.$$837_{\pm.009}.$ $1.410_{\pm.139}$$450_{\pm.075}.$ $2.141_{\pm.400}$
Jaccard .$761_{\pm.013}.$$850_{\pm.011}.$ $1.289_{\pm.181}$$502_{\pm.088}.$ $1.988_{\pm.641}$  .$766_{\pm.020}.$$854_{\pm.015}.$ $1.271_{\pm.145}$$521_{\pm.051}.$ $1.623_{\pm.313}$
Tversky .$752_{\pm.020}.$$843_{\pm.016}.$ $1.309_{\pm.086}$$517_{\pm.052}.$ $1.638_{\pm.369}$  .$768_{\pm.014}.$$856_{\pm.011}.$ $1.265_{\pm.129}$$521_{\pm.048}.$ $1.702_{\pm.298}$
WCE .$741_{\pm.013}.$$835_{\pm.011}.$ $1.417_{\pm.069}$$462_{\pm.019}.$ $2.100_{\pm.498}$  .$733_{\pm.009}.$$828_{\pm.007}.$ $1.441_{\pm.031}$$444_{\pm.039}.$ $2.326_{\pm.848}$
ABW -- -- -- -- --  -- -- -- -- --
Skea-topo -- -- -- -- --  .$599_{\pm.121}.$$674_{\pm.169}.$ $1.602_{\pm.113}$$265_{\pm.250}.$ $5.791_{\pm 4.492}$
RMI .$787_{\pm.012}.$$869_{\pm.010}.$ $1.282_{\pm.127}$$449_{\pm.064}.$ $1.797_{\pm.489}$  .$785_{\pm.009}.$$867_{\pm.007}.$ $1.279_{\pm.096}$$448_{\pm.047}.$ $1.884_{\pm.340}$
clDice .$482_{\pm.094}.$$598_{\pm.085}.$ $2.010_{\pm.119}$$207_{\pm.129}.$ $9.813_{\pm 8.153}$  .$509_{\pm.063}.$$617_{\pm.066}.$ $1.912_{\pm.157}$$238_{\pm.161}.$ $15.704_{\pm 11.080}$
Sensitive .$763_{\pm.008}.$$850_{\pm.007}.$ $1.440_{\pm.060}$$376_{\pm.039}.$ $2.824_{\pm.562}$  .$757_{\pm.018}.$$846_{\pm.015}.$ $1.444_{\pm.040}$$380_{\pm.028}.$ $3.134_{\pm.782}$
CWMI-Real .$\underline{.788_{\pm.016}}$$\underline{.870_{\pm.012}}$ $\underline{1.104_{\pm.175}}^{*}$$\underline{.582_{\pm.075}}$ $\underline{1.309_{\pm.299}}^{*}$  .$\underline{.787_{\pm.026}}$$\underline{.870_{\pm.019}}$ $\underline{1.072_{\pm.113}}^{**}$$\underline{.594_{\pm.057}}^{*}$ $\underline{1.398_{\pm.468}}$
CWMI .$\mathbf{.798_{\pm.012}}^{*}$$\mathbf{.878_{\pm.009}}^{*}$ $\mathbf{1.032_{\pm.176}}^{*}$$\mathbf{.613_{\pm.082}}^{*}$ $\mathbf{1.079_{\pm.382}}^{*}$  .$\mathbf{.795_{\pm.015}}^{*}$$\mathbf{.875_{\pm.012}}^{*}$ $\mathbf{1.007_{\pm.162}}^{**}$$\mathbf{.622_{\pm.062}}^{**}$ $\mathbf{1.219_{\pm.410}}^{*}$
@@
CE .$733_{\pm.015}.$$826_{\pm.013}.$ $3.473_{\pm.608}$$194_{\pm.083}.$ $1.927_{\pm 3.301}$  .$640_{\pm.209}.$$725_{\pm.195}.$$906_{\pm.107}.$$406_{\pm.327}.$ $7.914_{\pm 6.637}$
BCE .$689_{\pm.011}.$$794_{\pm.009}.$ $1.708_{\pm.265}$$560_{\pm.059}.$ $12.792_{\pm 3.684}$  .$619_{\pm.222}.$$708_{\pm.208}.$$986_{\pm.139}.$$365_{\pm.358}.$ $4.354_{\pm 1.869}$
Dice .$761_{\pm.011}.$$849_{\pm.009}.$ $1.601_{\pm.138}$$558_{\pm.042}.$ $11.189_{\pm 4.532}$  .$636_{\pm.197}.$$723_{\pm.187}.$$954_{\pm.056}.$$396_{\pm.322}.$ $7.335_{\pm 4.843}$
Focal .$671_{\pm.015}.$$780_{\pm.013}.$ $1.717_{\pm.150}$$556_{\pm.042}.$ $11.576_{\pm 1.950}$  .$581_{\pm.252}.$$672_{\pm.240}.$ $1.043_{\pm.109}$$336_{\pm.369}.$ $5.040_{\pm 2.802}$
Jaccard .$759_{\pm.008}.$$848_{\pm.006}.$ $1.332_{\pm.041}$$630_{\pm.008}.$ $11.274_{\pm 2.805}$  .$638_{\pm.202}.$$725_{\pm.190}.$$888_{\pm.076}.$$406_{\pm.331}.$ $6.758_{\pm 4.431}$
Tversky .$755_{\pm.009}.$$845_{\pm.007}.$ $1.379_{\pm.196}$$618_{\pm.061}.$ $11.501_{\pm 4.470}$  .$627_{\pm.195}.$$717_{\pm.186}.$$924_{\pm.102}.$$384_{\pm.314}.$ $7.521_{\pm 4.493}$
WCE .$650_{\pm.016}.$$762_{\pm.014}.$ $1.726_{\pm.103}$$556_{\pm.004}.$ $1.845_{\pm.470}$  .$823_{\pm.016}.$$898_{\pm.011}.$$873_{\pm.097}.$$705_{\pm.023}.$ $2.590_{\pm.602}$
ABW .$685_{\pm.014}.$$791_{\pm.012}.$ $1.678_{\pm.124}$$566_{\pm.030}.$ $11.654_{\pm 4.661}$  .$762_{\pm.019}.$$858_{\pm.013}.$ $1.472_{\pm.103}$$614_{\pm.037}.$ $\underline{3.623_{\pm.887}}$
Skea-topo .$620_{\pm.095}.$$734_{\pm.085}.$ $2.236_{\pm.824}$$463_{\pm.159}.$ $11.653_{\pm 5.620}$  .$784_{\pm.016}.$$873_{\pm.010}.$ $1.299_{\pm.086}$$645_{\pm.022}.$ $3.363_{\pm.609}$
RMI .$\underline{.766_{\pm.008}}$$852_{\pm.006}.$ $1.824_{\pm.263}$$499_{\pm.065}.$ $11.632_{\pm 3.918}$  .$839_{\pm.013}.$$908_{\pm.008}.$$819_{\pm.071}.$$727_{\pm.032}.$ $2.587_{\pm.435}$
clDice .$645_{\pm.051}.$$755_{\pm.045}.$ $2.162_{\pm.467}$$476_{\pm.081}.$ $15.461_{\pm 4.999}$  .$800_{\pm.013}.$$884_{\pm.009}.$ $1.004_{\pm.113}$$684_{\pm.026}.$ $3.146_{\pm.286}$
Sensitive .$734_{\pm.022}.$$827_{\pm.018}.$ $2.933_{\pm.982}$$305_{\pm.159}.$ $11.292_{\pm 3.828}$  .$824_{\pm.018}.$$899_{\pm.011}.$$880_{\pm.108}.$$702_{\pm.041}.$ $2.757_{\pm.398}$
CWMI-Real .$765_{\pm.010}.$$\underline{.853_{\pm.008}}$ $\underline{1.231_{\pm.183}}$$\underline{.650_{\pm.050}}$ $\underline{1.836_{\pm 4.994}}$  .$\underline{.846_{\pm.012}}$$\underline{.913_{\pm.008}}$$\underline{.766_{\pm.067}}^{**}$$\underline{.736_{\pm.021}}^{*}$ $2.504_{\pm.528}$
CWMI .$\mathbf{.767_{\pm.010}}$$\mathbf{.854_{\pm.008}}$ $\mathbf{1.148_{\pm.166}}^{*}$$\mathbf{.667_{\pm.047}}^{*}$ $\mathbf{1.326_{\pm 4.109}}$  .$\mathbf{.839_{\pm.015}}$$\mathbf{.908_{\pm.010}}$$\mathbf{.762_{\pm.084}}^{***}$$\mathbf{.728_{\pm.035}}^{**}$ $\mathbf{2.747_{\pm.680}}$
[…]
@@ 4.1 Experimental Setup
To evaluate the effectiveness of the CWMI loss, we employed U-Net (Ronneberger et al., 2015) and Attention U-Net (Oktay et al., 2018) as baseline architectures. This selection enables an assessment of CWMI’s generalization ability across both fully convolutional and attention-enhanced models. Additionally, to examine CWMI’s compatibility with recently proposed Mamba-based architectures (Yue & Li, 2024), we incorporated the Vision Mamba U-Net (VMUNet) (Ruan et al., 2024) into our experiments.
@@
We tested CWMI on three public segmentation datasets, all characterized by class and instance imbalance: (1) SNEMI3D, a neurite segmentation dataset containing 100 $1024\times 1024$ grayscale images from electron microscopy slices (Arganda-Carreras et al., 2013); (2) GlaS, a gland segmentation dataset with 165 RGB images of varying sizes from histological images of colorectal cancer samples (Sirinukunwattana et al., 2017); (3) DRIVE, a retinal vessel segmentation dataset comprising 40 $584\times 565$ RGB images from fundus photographs (Staal et al., 2004); and (4) the Massachusetts Roads dataset (MASS ROAD), a road segmentation dataset with 1171 $1500\times 1500$ RGB images from aerial imagery (Mnih, 2013). We choose a subset of 120 images (ignoring images without a network of roads). For all datasets, at least two repeated three-fold cross-validations were used to ensure robust evaluation.
@@
[…]
We utilized steerable pyramids with four decomposition levels and four orientations in all experiments. $N=4$ and $K=4$ of the steerable pyramid decomposition and a regularization parameter of $\lambda=0.1$ were determined by the ablation experiments, and applied in all CWMI experiments. For the implementation of Equation 9, although PyTorch supports complex matrix calculations, our experiments indicated that its efficiency remains suboptimal. Consequently, we computed using their real representations, which are mathematically equivalent (Golub & Van Loan, 2013). To assess the significance of the complex steerable pyramid, we compared CWMI with a real-number-only variant (CWMI-Real), implemented according to Equations 4 and 8. Adam optimizer with a StepLR scheduler (initial learning rate $1\times 10^{-4}$, decay rate 0.8, step size 10) was used. U-Net and Attention U-Net models were trained for 50 epochs; VMUNet models were trained for 100 epochs due to their slower convergence. All models were trained with a batch size of 10. Early stopping based on mIoU was employed to select the best model. Training was conducted on an NVIDIA A100 GPU using the Google Colab runtime.
@@
     SNEMI3D VMUNet
[…]
CE .$723_{\pm.026}.$$829_{\pm.020}.$ $2.577_{\pm 1.211}$$414_{\pm.173}.$ $1.709_{\pm.235}$
BCE .$739_{\pm.004}.$$843_{\pm.002}.$ $1.499_{\pm.057}$$\underline{.594_{\pm.010}}$ $1.274_{\pm.231}$
Dice .$725_{\pm.029}.$$831_{\pm.021}.$ $2.735_{\pm 1.213}$$402_{\pm.165}.$ $1.767_{\pm.720}$
Focal .$725_{\pm.010}.$$833_{\pm.007}.$ $1.677_{\pm.141}$$563_{\pm.025}.$ $1.699_{\pm.372}$
Jaccard .$725_{\pm.019}.$$831_{\pm.014}.$ $1.953_{\pm.541}$$499_{\pm.075}.$ $1.716_{\pm.566}$
Tversky .$737_{\pm.016}.$$840_{\pm.012}.$ $1.619_{\pm.324}$$555_{\pm.045}.$ $1.560_{\pm.408}$
WCE .$714_{\pm.005}.$$826_{\pm.004}.$ $1.743_{\pm.026}$$552_{\pm.006}.$ $1.432_{\pm.169}$
ABW .$678_{\pm.003}.$$800_{\pm.003}.$ $2.076_{\pm.030}$$490_{\pm.005}.$ $1.572_{\pm.021}$
RMI .$\underline{.762_{\pm.010}}$$\underline{.857_{\pm.007}}$ $\underline{1.419_{\pm.183}}$$588_{\pm.030}.$ $\underline{1.071_{\pm.306}}$
clDice .$714_{\pm.002}.$$824_{\pm.001}.$ $1.754_{\pm.062}$$542_{\pm.003}.$ $1.990_{\pm.262}$
sensitive -- -- -- -- --
CWMI .$\mathbf{.783_{\pm.004}}^{*}$$\mathbf{.872_{\pm.003}}^{*}$$\mathbf{.982_{\pm.106}}^{*}$$\mathbf{.660_{\pm.022}}^{*}$$\mathbf{.629_{\pm.073}}$
[…]
Figure 3: Qualitative results of different loss functions on the SNEMI3D dataset. Red: false positive regions; Blue: false negative regions. Green arrow: challenging false positive and Orange arrow: challenging false negative that are successfully addressed by CWMI.